OpenStudy (anonymous):

if the volume of a sphere above is 4,500 π cubic units , what is the radius of the sphere A.) 15units B.)60units C.)5units D.)30units

OpenStudy (anonymous):

katie, lets do this together.

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

do you know the formula for the volume of a sphere with radius r ?

OpenStudy (anonymous):

Ummm no i don't sorry

OpenStudy (anonymous):

do i click on the link?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

okay i got it. Thanks so thats step one

OpenStudy (anonymous):

okay great. so 4/3 pi r^3 = 4500 pi

OpenStudy (anonymous):

okay yes :)

OpenStudy (anonymous):

so solve for r.

OpenStudy (anonymous):

do i times a 4500 times 3?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

and then?

OpenStudy (anonymous):

okay i got 13500 :)

OpenStudy (anonymous):

what does 13500 equal to?

OpenStudy (anonymous):

15 units

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

sorry i get the first part but i don't know what you mean whats it equal to

OpenStudy (anonymous):

okay lets start over. \[4/3 * \pi * r ^{3} = 4500 \pi\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so, \[r ^{3} = 4500 \pi / (\pi*4/3) = 4500*3/4 \]

OpenStudy (anonymous):

okay which is 3375 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and so the radius is?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

i looke at the formula

OpenStudy (anonymous):

okay and?

OpenStudy (anonymous):

idk is the radus 3375

OpenStudy (anonymous):

no. \[r ^{3} = 3375 \rightarrow r = ?\]

OpenStudy (anonymous):

if x^3 is 8, what is x?

OpenStudy (anonymous):

8x3=24 so 24

OpenStudy (anonymous):

okay x^3 means x cubed or x raised to the power 3. x is 2. 2 cube is 8. 2 raised to the power 3 is 8. In other words, 2 multiplied by itself 2 times is 8. i.e 8 = 2 times 2 times 2

OpenStudy (anonymous):

sorry multiplied by itself 3 times.

OpenStudy (anonymous):

2 is called the cube-root of 8.

OpenStudy (anonymous):

similarly radius cubed is 3375. so radius is cube root of 3375.

OpenStudy (anonymous):

use your calculator to find the cube root of 3375. that is your radius.

OpenStudy (anonymous):

okay which is 7425

OpenStudy (anonymous):

do you know what square root is?

OpenStudy (anonymous):

yes hold on :0

OpenStudy (anonymous):

okay 61.4410

OpenStudy (anonymous):

what is 61.441 ?

OpenStudy (anonymous):

the square root

OpenStudy (anonymous):

of?

OpenStudy (anonymous):

3375

OpenStudy (anonymous):

sqrt(3 375) = 58.0947502

OpenStudy (anonymous):

I was asking you if you know the concept of square root.

OpenStudy (anonymous):

ohhh so the answer is b 60 units

OpenStudy (anonymous):

NO. pay attention.

OpenStudy (anonymous):

leave the problem for a moment. Do you know the concept of square root?

OpenStudy (anonymous):

i guess not..

OpenStudy (anonymous):

okay. if you multiply a number by itself. it will give you another number. For example, if you multiply 4 by itself, that is, multiply 4 by 4, you get 16.

OpenStudy (anonymous):

16 is the square of 4.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

and 4 is the square root of 16.

OpenStudy (anonymous):

do you know the area of a square?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

area of a square of side A is A*A. that is why this is called square and square root. anyway, thats not important,

OpenStudy (anonymous):

similarly, if you multiply a number by itself 3 times, you get another number. so if you multiply 4 by itself thrice, that is 4 times 4 times 4, you get 64.

OpenStudy (anonymous):

64 is the cube of 4. conversely, 4 is the cube root of 64.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

okay so, if the radius is multiplied by itself 3 times, you get 3375. so, radius*radius*radius = 3375

OpenStudy (anonymous):

so, to get radius, we need to find the cube root of 3375.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

now, 10*10*10 = 1000

OpenStudy (anonymous):

and 20*20*20 = 8000

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so r*r*r = 3375. 3375 lies between 1000 and 8000. so r must lie between 10 and 20. So, r is 15 units.

OpenStudy (anonymous):

ohhhhhh okay thank you so much :)

OpenStudy (anonymous):

you are welcome.